# Bitcoin Under Volatile Block Rewards: How Mempool Statistics Can Influence Bitcoin Mining

### v1

@@ 5. Enhanced Volatile Reward Model
This section introduces our Asynchronous Advantage Actor Critic (A3C)-based implementation designed to derive a near-optimal mining policy within a more realistic volatile reward model. A brief discussion of the A3C algorithm is provided in Appendix B. Our implementation codes are presented in (our, 2024). The environment used by the A3C-based tool includes a transaction mempool, allowing us to realistically simulate the behavior of the Bitcoin mempool. Our A3C-based implementation can analyze mining attacks such as selfish mining and undercutting, both before and after the difficulty adjustment mechanism.
As discussed in Appendix B, A3C is a reinforcement learning algorithm where the agent (the payoff-maximizing miner) interacts with the environment (the Bitcoin blockchain). Through these interactions, the agent transitions between different states, each containing the necessary information to represent both blockchain forks and mempool conditions. At each state, the adversary can choose from various possible actions, with each action offering different rewards. A detailed discussion of the possible action set and state representation in our implementation is presented in Appendix E.1.
@@ 5.1. Enhanced Mempool Implementation
In the simplified model, we estimate the time-fee equation using only data on block transaction fees and generation times. However, the way transaction fees impact block rewards is more complex in practice. To better estimate how transaction fees fluctuate over time, we can improve the model by incorporating the behavior of the Bitcoin mempool. The mempool is a storage space where unconfirmed transactions are stored before they are included in a block. To increase the chance of inclusion in a block, each Bitcoin transaction pays a fee to the block miner. The total fee that a transaction pays is calculated based on the transaction’s weight, which is measured in virtual bytes (vBytes). Virtual bytes are a weight unit introduced by the SegWit upgrade. The fee is typically expressed in satoshis per virtual byte (sat/vByte), where each satoshi is $10^{-8}$ BTC. Each Bitcoin block has a size limit of 1 virtual megabytes (vMB) (Bit, 2024). Miners prioritize filling the block with the most valuable transactions from the mempool, where the term valuable refers to transactions that offer the highest fees in satoshis per virtual byte. In the context of the enhanced volatile reward model, we define the base fee for a given mempool period as the fee for which there is at least 1 virtual megabyte of transactions offering that fee or higher at all times during that period.
Our implementation categorizes the transactions available in the mempool into $N^{\text{range}}$ sat/vByte ranges, denoted as
$\{\texttt{fee}_{0},\texttt{fee}_{1},\ldots,\texttt{fee}_{N^{\text{range}}-1}\}\enspace,$
@@
where the fee levels are ordered in ascending sequence. A transaction paying $f$ sat/vByte, where $\texttt{fee}_{j}\leq f<\texttt{fee}_{j+1}$, is assigned to the range corresponding to $\texttt{fee}_{j}$. $\texttt{fee}_{0}$ denotes the base fee, implying that all transactions available in the mempool pay at least $\texttt{fee}_{0}$ sat/vByte. In our implementation, we assume an unlimited number of transactions paying the base fee. The weight of transactions in the range $\texttt{fee}_{j}$ increases over time according to a specific equation, called the weight-time equation, which is estimated based on the transaction weight growth in the range $\texttt{fee}_{j}$. Whenever a new block is mined, a total weight equivalent to 1 virtual megabyte is deducted from mempool, starting from the highest sat/vByte ranges. To estimate the weight increase function for each sat/vByte range over a given time period, we analyzed mempool statistics during that period using the information presented in (mem, 2024). Using these statistics, we applied regression techniques to derive the weight-time equation for each sat/vByte range.
5.2. Objective Function
5.2.1. Before the difficulty adjustment
5.2.2. After the difficulty adjustment
Since the objective function defined in equation 9 is non-linear, we cannot naively assign the time step reward as is done in Section 5.2.1. To address this challenge, we adopt the approach introduced in (Sapirshtein et al., 2017), which converts the non-linear optimization function into a linear one. We set the time step reward as follows:
where $\rho$ represents the time-averaged profit the adversary earns by following the optimal strategy, i.e., $\rho=O^{\mathcal{A}}(\pi^{*})$, where $\pi^{*}$ denotes the optimal strategy. At the start of training, the actual value of $\rho$ is unknown. Therefore, the training agents initialize with $\rho=O^{\mathcal{A}}(\pi^{\mathcal{H}})$, where $\pi^{\mathcal{H}}$ denotes the honest strategy. Certain test agents are tasked with measuring the time-averaged profit under the policy trained by the training agents at specific time steps during the training process. If this profit exceeds the highest time-averaged profit achieved so far, they prompt the training agents to update $\rho$ to the new highest time-averaged profit. For the detailed analysis of the non-linear objective function conversion readers are referred to (Sapirshtein et al., 2017). We briefly provide the intuition behind setting the time step reward as defined in equation 10 in Appendix F.1.
Another critical point is that the average time step duration after difficulty adjustment needs recalibration in the environment. For details on how our implementation measures the time step duration, readers are referred to Appendix F.2.
@@ 5.3. Results and Implications: Mining Attacks Under an Enhanced Volatile Model
This section presents results from our A3C-based implementation (our, 2024), analyzing mining attacks under the mempool patterns observed during different periods. Information regarding our A3C-based tool implementation is provided in E.4. To compare the time-averaged profits of a mining strategy $\pi$, we use the concept of the percentage increase in the miner’s time-averaged profit, defined as follows:
@@
In this section, we analyze the profitability of selfish mining before the difficulty adjustment, as well as the single-block undercutting attack. The results related to the profitability of selfish mining after the difficulty adjustment is presented in Appendix G.
@@
Similarly, Figure 8 illustrates the percentage increase in the adversary’s time-averaged profit before a difficulty adjustment, using the same assumptions as in Figure 7, with the exception that the protocol reward is set to zero. By comparing Figures 8 and 7, we observe that as the protocol reward decreases, deviations from the honest strategy become increasingly profitable, posing a potential threat to Bitcoin’s progress in the near future.
@@
Although Figure 8 shows that in the transaction-fee era, the same mempool pattern as occurred on December 18, 2023, can be concerning, there are some other days where the threat of adversarial attack is more alarming. Figure 9 represents the percentage increase in profit before the difficulty adjustment from applying both selfish mining and undercutting attacks simultaneously under the mempool patterns of two different periods: December 18, 2023, between 23:00 and 24:00 CET (the same period as in Figures 8 and 7), and October 15, 2023, between 8:00 and 9:00 CET. The figure is depicted as a function of different ratios of petty-compliant miners under the following assumptions: i) the protocol reward equals zero, ii) the adversarial mining share is $\frac{1}{3}$, and iii) the adversary has a normal communication capability of 50%. As can be seen in Figure 9, even if all the honest miners are altruistic, the deviation still yields a substantial immediate profit for the adversary under the mempool pattern of October 15, 2023. By increasing the ratio of petty-compliant miners in the system, the deviation can become even more profitable. The main reason for the higher profitability of mining attacks under the mempool pattern of October 15, 2023, compared to that of December 18, 2023, is the low level of base fee available in the mempool. Between 8:00 and 9:00 CET on October 15, 2023, the base fee was 1 sat/vByte (mem, 2024), implying that during this period, miners needed to fill at least some part of their blocks with transactions offering only 1 sat/vByte. In contrast, the base fee was 80 sat/vByte on December 18, 2023, between 23:00 and 24:00 CET (mem, 2024). This comparison shows that a higher base fee level for transactions available in the mempool can help mitigate mining attacks in the transaction-fee era.
@@
In our A3C-based implementation, we treat the undercut action as a continuous action, where the agent must not only decide whether to take this action but also determine the duration for which it should perform undercutting. Once the designated time has passed, the agent should stop undercutting the tip block and begin mining on top of it, as a sufficient number of transactions will have arrived in the system. For more information on the continuity of the undercut action, readers are referred to Appendix E.2.
To measure the profitability of the single-block undercutting attack, we have limited the adversary’s possible action set, as outlined in Appendix E.3. Figure 10 shows the percentage increase in time-averaged profit achieved by performing the undercutting attack. The figure is depicted based on mempool patterns observed over five different days between 8:00 and 9:00 CET, under the following assumptions: i) the protocol reward is set to 0, and ii) the adversary’s communication capability equals 1.
@@ Appendix B Asynchronous Advantage Actor-Critic (A3C)
Appendix C Proof of Theorem 3.1

### v2

@@ 6. Enhanced Volatile Reward Model
This section introduces our Asynchronous Advantage Actor Critic (A3C)-based implementation designed to derive a near-optimal mining policy within a more realistic volatile reward model. Both the WeRLman model (Bar-Zur et al., 2023) and our simplified model introduced in Section 4 rely on assumptions that overlook the complexity of the real-world mempool. For instance, both models simplify transaction fee rewards as discrete, and the simplified model assumes linear time-fee equations, whereas in practice, time-fee relationships are typically concave. Moreover, the simplified model assumes that a block’s fee reward depends solely on its generation time, disregarding the correlation between a block’s transaction fee reward and the generation times of its preceding blocks. We elaborate further on the limitations of the simplified model in the Appendix E. These limitations highlight the need for a more nuanced approach that incorporates the behavior of the Bitcoin mempool to provide a more precise analysis of mining strategies in the transaction-fee era. To address these challenges, we introduce an enhanced mempool implementation, designed to mimic real-world mempool behavior within the environment. Given the large state space of this enhanced environment, traditional MDP tools are no longer viable. Instead, we leverage A3C reinforcement learning techniques to explore and analyze mining strategies effectively. Our A3C-based implementation, which considers a wider variety of actions, can analyze adversarial strategies such as selfish mining and undercutting, both before and after the difficulty adjustment mechanism.
Our A3C-based implementation codes are provided in (our, 2024). A brief introduction to the A3C algorithm is presented in Appendix B. Additionally, a detailed discussion on the implementation of the A3C-based tool is provided in Appendix F. This discussion includes an explanation of the possible action set and state representation in our implementation (Appendix F.1), a comprehensive description of the objective function (Appendix F.2), and the implementation details of the undercutting action as a continuous action (Appendices F.3 and F.4).
@@ 6.1. Enhanced Mempool Implementation
In the simplified model, we estimate the time-fee equation using only data on block transaction fees and generation times. To better estimate how transaction fees fluctuate over time, we can improve the model by incorporating the behavior of the Bitcoin mempool. The mempool is a storage space where unconfirmed transactions are stored before they are included in a block. To increase the chance of inclusion in a block, each Bitcoin transaction pays a fee to the block miner. The total fee that a transaction pays is calculated based on the transaction’s weight, which is measured in virtual bytes (vBytes). Virtual bytes are a weight unit introduced by the SegWit upgrade. The fee is typically expressed in satoshis per virtual byte (sat/vByte), where each satoshi is $10^{-8}$ BTC. Each Bitcoin block has a size limit of 1 virtual megabytes (vMB) (Bit, 2024). Miners prioritize filling the block with the most valuable transactions from the mempool, where the term valuable refers to transactions that offer the highest fees in satoshis per virtual byte. In the context of the enhanced volatile reward model, we define the base fee for a given mempool period as the fee for which there is at least 1 virtual megabyte of transactions offering that fee or higher at all times during that period.
@@
Our implementation categorizes the transactions available in the mempool into $N^{\text{range}}$ sat/vByte ranges, denoted as $\{\texttt{fee}_{0},\texttt{fee}_{1},\ldots,$ $\texttt{fee}_{N^{\text{range}}-1}\}\enspace,$ where the fee levels are ordered in ascending sequence. A transaction paying $f$ sat/vByte, where $\texttt{fee}_{j}\leq f<\texttt{fee}_{j+1}$, is assigned to the range corresponding to $\texttt{fee}_{j}$. $\texttt{fee}_{0}$ denotes the base fee, implying that all transactions available in the mempool pay at least $\texttt{fee}_{0}$ sat/vByte. In our implementation, we assume an unlimited number of transactions paying the base fee. The weight of transactions in the range $\texttt{fee}_{j}$ increases over time according to a specific equation, called the weight-time equation, which is estimated based on the transaction weight growth in the range $\texttt{fee}_{j}$. Whenever a new block is mined, a total weight equivalent to 1 virtual megabyte is deducted from mempool, starting from the highest sat/vByte ranges. To estimate the weight increase function for each sat/vByte range over a given time period, we analyzed mempool statistics during that period using the information presented in (mem, 2024). Then, we applied regression techniques to derive the weight-time equation for each sat/vByte range. The estimated result of a weight-time equation for a sample day is presented in Appendix G.
@@ 6.2. Results and Implications: Mining Attacks Under an Enhanced Volatile Model
This section presents results from our A3C-based implementation (our, 2024), analyzing mining attacks under the mempool patterns observed during different periods. Information regarding our A3C-based tool implementation is provided in F.5. To compare the time-averaged profits of a mining strategy $\pi$, we use the concept of the percentage increase in the miner’s time-averaged profit, defined as follows:
@@
In this section, we analyze the profitability of selfish mining before the difficulty adjustment, as well as the single-block undercutting attack. The results related to the profitability of selfish mining after the difficulty adjustment is presented in Appendix H.
@@
Similarly, Figure 10 illustrates the percentage increase in the adversary’s time-averaged profit before a difficulty adjustment, using the same assumptions as in Figure 9, with the exception that the protocol reward is set to zero. By comparing Figures 10 and 9, we observe that as the protocol reward decreases, deviations from the honest strategy become increasingly profitable, posing a potential threat to Bitcoin’s progress in the near future.
@@
Although Figure 10 shows that in the transaction-fee era, the same mempool pattern as occurred on December 18, 2023, can be concerning, there are some other days where the threat of adversarial attack is more alarming. Figure 11 represents the percentage increase in profit before the difficulty adjustment from applying both selfish mining and undercutting attacks simultaneously under the mempool patterns of two different periods: December 18, 2023, between 23:00 and 24:00 CET (the same period as in Figures 10 and 9), and October 15, 2023, between 8:00 and 9:00 CET. The figure is depicted as a function of different ratios of petty-compliant miners under the following assumptions: i) the protocol reward equals zero, ii) the adversarial mining share is $\frac{1}{3}$, and iii) the adversary has a normal communication capability of 50%. As can be seen in Figure 11, even if all the honest miners are altruistic, the deviation still yields a substantial immediate profit for the adversary under the mempool pattern of October 15, 2023. By increasing the ratio of petty-compliant miners in the system, the deviation can become even more profitable. The main reason for the higher profitability of mining attacks under the mempool pattern of October 15, 2023, compared to that of December 18, 2023, is the low level of base fee available in the mempool. Between 8:00 and 9:00 CET on October 15, 2023, the base fee was 1 sat/vByte (mem, 2024), implying that during this period, miners needed to fill at least some part of their blocks with transactions offering only 1 sat/vByte. In contrast, the base fee was 80 sat/vByte on December 18, 2023, between 23:00 and 24:00 CET (mem, 2024). This comparison shows that a higher base fee level for transactions available in the mempool can help mitigate mining attacks in the transaction-fee era.
@@
In our A3C-based implementation, we treat the undercut action as a continuous action, where the agent must not only decide whether to take this action but also determine the duration for which it should perform undercutting. Once the designated time has passed, the agent should stop undercutting the tip block and begin mining on top of it, as a sufficient number of transactions will have arrived in the system. For more information on the continuity of the undercut action, readers are referred to Appendix F.3.
To measure the profitability of the single-block undercutting attack, we have limited the adversary’s possible action set, as outlined in Appendix F.4. Figure 12 shows the percentage increase in time-averaged profit achieved by performing the undercutting attack. The figure is depicted based on mempool patterns observed over five different days between 8:00 and 9:00 CET, under the following assumptions: i) the protocol reward is set to 0, and ii) the adversary’s communication capability equals 1.
@@ Appendix B Asynchronous Advantage Actor-Critic (A3C)
B.1. Why Use A3C for Selfish Mining Strategy
As mentioned in (Han et al., 2023), actor–critic algorithms (Konda and Tsitsiklis, 1999) combine value-based and policy-based reinforcement learning (RL) methods. The actor, a policy network, proposes actions for given states, while the critic, a value network, evaluates these actions based on the state-action pair. By leveraging the Bellman equation, the critic updates the Q-function, which in turn guides the actor’s policy updates, merging the strengths of both RL paradigms.
The Asynchronous Advantage Actor–Critic (A3C) algorithm (Mnih et al., 2016), introduced by DeepMind, extends this framework by employing asynchronous parallelism. Multiple agents interact independently with different environment copies, enabling faster exploration and improved coverage of the state-action space. This makes A3C particularly efficient for tasks involving dynamic and high-dimensional environments.
A3C is especially well-suited for optimizing strategies for selfish mining in a volatile reward model. The state complexity significantly increases in our investigation of selfish mining within a simulated mining pool. Consequently, we require a method capable of handling high-dimensional environments, and A3C meets this need. Additionally, using multiple agents in A3C boosts the speed of exploration and learning, further enhancing its efficiency in these complex scenarios.
Appendix C Non-Predictable WeRLman
In order to modify the original WeRLman environment to a non-predictable version, we updated the function
tryGetWithAndWithoutNewFee in the file
bitcoin_simplified_fee_model.py, which is available in the repository referenced in (Wer, 2022), as presented in Algorithm 1.
Algorithm 1 Updated function tryGetWithAndWithoutNewFee
1:Input: Previous state element prev_state.
2:Output: Updated states new_with_fee, new_without_fee.
3:$current\_max\_fork\leftarrow\max(h,a)$
[…]
5:Create a copy of the current state: $new\_with\_fee\leftarrow\text{copy}()$
6:Create another copy of the current state: $new\_without\_fee\leftarrow\text{copy}()$
[…]
8:     Increment the pool of new_with_fee: $new\_with\_fee.pool\leftarrow new\_with\_fee.pool+1$
[…]
10:     if $new\_honest$ then
[…]
12:         Increment T_h of new_with_fee: $new\_with\_fee.T\_h\leftarrow new\_with\_fee.T\_h+with\_fee$
13:     else
14:         if $a\leq h$ then
[…]
16:              Increment T_a of new_with_fee: $new\_with\_fee.T\_a\leftarrow new\_with\_fee.T\_a+with\_fee$
17:         else
18:              Set the last element of L: $new\_with\_fee.L[-1]\leftarrow 1$
19:         end if
20:     end if
21:end if
22:Return new_with_fee, new_without_fee
Appendix D Adversarial Strategies in the Original and Non-predictable WeRLman Environments
Let $n_{\mathcal{A}}$, $n_{\mathcal{H}}$ denote the number of blocks in the adversarial and the honest fork, respectively. Let $\alpha$, $g$, $P$, and $F$ denote the adversarial mining share, the adversarial communication capability, the whale transaction probability, and the extra fee of a whale transaction, respectively. We denote by $N_{\mathcal{A}}$, $N_{\mathcal{H}}$, and $R_{\mathcal{A}}$ the normalized average number of canonical adversarial blocks, the normalized average number of canonical honest blocks, and the normalized average adversarial block reward. The time-averaged profit of strategy $\pi$ can be calculated as $\texttt{Profit}(\pi)=\frac{R_{\mathcal{A}}{(\pi)}}{N_{\mathcal{A}}{(\pi)}+N_{\mathcal{H}}{(\pi)}}$. Note that the time-averaged profit of the adversary if mining honestly is equal to $\texttt{Profit}(\pi^{\mathcal{H}})=\alpha(1_{F}P)$.
D.1. Strategy $\pi_{1}^{\text{WeRLman}}$
Let $\texttt{tnx}=0$ denote the existence of a normal transaction in the pool, and $\texttt{tnx}=1$ denote the existence of a whale transaction in the pool. The state space and the action in each state are as follows:
State $S_{0}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0,\texttt{tnx}=0)$: If the next block is honest: adopt. If the next block is adversarial and the next transaction is a whale: wait. If the next block is adversarial and the next transaction is normal: publish.
State $S_{0}^{\prime}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0,\texttt{tnx}=1)$: If the next block is honest: adopt. If the next block is adversarial: publish.
State $S_{1}:(n_{\mathcal{A}}=1(\texttt{normal}),n_{\mathcal{H}}=0,\texttt{tnx}=1)$: If the next block is adversarial: wait until $n_{\mathcal{A}}-n_{\mathcal{H}}=1$, then Override. If the next block is honest: match.
State $S_{2}:(n_{\mathcal{A}}=1(\texttt{normal-published}),n_{\mathcal{H}}=1)$: In this state, a new transaction is not sampled, as the same height block already exists. If the next block is honest: adopt. If the next block is adversarial: publish.
Let $P_{0}$, $P_{0}^{\prime}$, $P_{1}$, $P_{2}$ denote the stationary probability of being in states $S_{0}$, $S_{0}^{\prime}$, $S_{1}$, $S_{2}$, respectively. We have:
(11) $\begin{split}&P_{0}=\frac{1-p}{1+p(1-p)\alpha(1-\alpha)},\quad P_{1}=\frac{\alpha p(1-p)}{1+p(1-p)\alpha(1-\alpha)},\quad\\ &P_{2}=\frac{p(1-p)\alpha(1-\alpha)}{1+p(1-p)\alpha(1-\alpha)},\quad P_{0}^{\prime}=1-P_{0}-P_{1}-P_{2}.\end{split}$
The normalized average number of canonical honest blocks, canonical adversarial blocks, can be obtained as follows:
(12) $\begin{split}N_{\mathcal{H}}=&P_{0}\left(1-\alpha\right)+2P_{2}\left(1-\alpha\right)\left(1-g\right)+P_{2}\left(1-\alpha\right)g+P_{0}^{\prime}\left(1-\alpha\right),\\ N_{\mathcal{A}}=&P_{0}\alpha\left(1-p\right)+\left(2+\frac{\alpha}{1-2\alpha}\right)P_{1}\alpha+2P_{2}\alpha+P_{2}\left(1-\alpha\right)g+P_{0}^{\prime}\alpha,\\ R_{\mathcal{A}}=&P_{0}\alpha\left(1-p\right)+P_{0}^{\prime}\alpha\left(1+F\right)+P_{2}\alpha\left(2+F\right)+P_{2}\left(1-\alpha\right)g\\ &+P_{1}\alpha\left(2+F\right)+P_{1}\alpha\frac{\alpha}{1-2\alpha}\left(1+pF\right)\enspace.\end{split}$
D.2. Strategy $\pi_{1}^{\text{Non-predictable}}$
Let $\texttt{tnx}=0$ denote if the transaction included in the adversarial block is a normal one. The state space and the action in each state are as follows:
State $S_{0}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0)$: If the next block is honest: adopt. If the next block is adversarial and includes a whale: publish. If the next block is adversarial and includes a normal transaction: wait.
State $S_{1}:(n_{\mathcal{A}}=1(\texttt{normal}),n_{\mathcal{H}}=0,\texttt{tnx}=0)$: If the next block is honest: match. If the next block is adversarial: publish.
State $S_{2}:(n_{\mathcal{A}}=1(\texttt{normal-published}),n_{\mathcal{H}}=1,\texttt{tnx}=0)$: If the next block is honest: adopt. If the next block is adversarial: publish.
Let $P_{0}$, $P_{1}$, $P_{2}$ denote the stationary probability of being in states $S_{0}$, $S_{1}$, and $S_{2}$, respectively. We have:
(13) $\begin{split}&P_{0}=\frac{1}{1+\left(1-p\right)\alpha\left(2-\alpha\right)},\quad P_{1}=\alpha\left(1-p\right)P_{0},\quad\\ &P_{2}=\left(1-\alpha\right)P_{1}.\end{split}$
The normalized average number of canonical honest blocks, canonical adversarial blocks, can be obtained as follows:
(14) $\begin{split}N_{\mathcal{H}}=&P_{0}\left(1-\alpha\right)+P_{2}\left(1-\alpha\right)g+2P_{2}\left(1-\alpha\right)\left(1-g\right),\\ N_{\mathcal{A}}=&P_{0}\alpha p+P_{1}\alpha\left(2+\frac{\alpha}{1-2\alpha}\right)+2P_{2}\alpha+P_{2}\left(1-\alpha\right)g,\\ R_{\mathcal{A}}=&P_{0}\alpha p\left(1+F\right)+P_{1}\alpha\left(1+\left(1+pF\right)\left(1+\frac{\alpha}{1-2\alpha}\right)\right)\\ &+P_{2}\left(1-\alpha\right)g+P_{2}\alpha\left(2+pF\right)\enspace.\end{split}$
D.3. Strategy $\pi_{2}^{\text{Non-predictable}}$
Let $\texttt{tnx}=1$ denote if the transaction included in the honest block is a whale transaction. The state space and the action in each state are as follows:
State $S_{0}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=0)$: If the next block is honest and includes a normal transaction: adopt. If the next block is honest and includes a whale transaction: wait. If the next block is adversarial: publish.
State $S_{1}:(n_{\mathcal{A}}=0,n_{\mathcal{H}}=1,\texttt{tnx}=1)$: If the next block is honest and includes a normal transaction: adopt. If the next block is honest and includes a whale transaction: adopt the first honest block. If the next block is adversarial: wait.
State $S_{2}:(n_{\mathcal{A}}=1(\texttt{whale-unpublished}),n_{\mathcal{H}}=1)$: If the next block is honest and includes a normal transaction: adopt. If the next block is honest and includes a whale transaction: adopt the first honest block. If the next block is adversarial: publish.
Let $P_{0}$, $P_{1}$, $P_{2}$ denote the stationary probability of being in states $S_{0}$, $S_{1}$, and $S_{2}$, respectively. We have:
(15) $\begin{split}&P_{0}=1-\left(1-\alpha^{2}\right)p,\quad P_{1}=\left(1-\alpha\right)p,\quad\\ &P_{2}=\alpha\left(1-\alpha\right)p.\end{split}$
The normalized average number of canonical honest blocks, canonical adversarial blocks, can be obtained as follows:
(16) $\begin{split}N_{\mathcal{H}}=&P_{0}\left(1-\alpha\right)\left(1-p\right)+P_{1}\left(1-\alpha\right)p+2P_{1}\left(1-\alpha\right)\left(1-p\right)+\\ &P_{2}\left(1-\alpha\right)p+2P_{2}\left(1-\alpha\right)\left(1-p\right),\\ N_{\mathcal{A}}=&P_{0}\alpha+2P_{2}\alpha,\\ R_{\mathcal{A}}=&P_{0}\alpha\left(1+Fp\right)+P_{2}\alpha\left(2+F(1+p)\right)\enspace.\end{split}$
Appendix E Limitations of the Simplified Environment
Non-linearity of the time-fee equation: In Fig. 4 presented in Section 4.1, a linear regression model is used to represent the time-fee equation. However, the block transaction fee does not necessarily increase linearly with the block generation time. This occurs when the rate of fee increase is not constant and decreases as the block generation time increases. In such cases, using alternative regression techniques rather than a linear model can provide a time-fee equation that better fits the behavior observed in the Bitcoin mempool. In Fig. 13(a), we compare linear regression with a curve regression model given by $f(t)=0.6414\;\texttt{ln}(t+6.5209)-0.8419$ to derive the time-fee equation for Bitcoin blocks mined on May 3, 2023. The coefficient of determination (R-squared) for the curve regression is $0.6243$, while for the linear regression it is $0.5180$. This indicates that the curve regression provides a better fit compared to the linear regression, suggesting that the fee increase rate diminishes with longer block generation times.
Dependence of fee rewards on the generation times of preceding blocks: Another technique that can lead to a more realistic mempool implementation is using multivariate regression instead of univariate regression. In univariate regression, the block transaction fee is estimated based solely on a single predictor, namely the block generation time. In contrast, multivariate regression allows us to account for additional predictors, such as the generation time of the parent block, in determining the block transaction fee. In Fig. 13(b), we depict the regression plane for the block transaction fee on March 18, 2024, based on two predictors: the block generation time and the parent block generation time. The resulting time-fee equation is $f(t_{1},t_{2})=0.1697+0.0079*t_{1}+0.0046*t_{2}$, where $t_{1}$ and $t_{2}$ represent the block generation time and the parent block generation time, respectively. Note that if the parent block generation time is longer, more fee-valuable transactions will likely remain in the mempool after the parent block is mined, allowing the next block to collect these transactions. As the time-fee equation shows, the fee increase rate with respect to the block generation time is higher than the fee increase rate with respect to the parent block generation time. This implies that the block generation time is a more effective predictor of the transaction fee compared to the parent block generation time.
(a) May 3, 2023
(b) March 18, 2024
Figure 13. Time-fee equation
Appendix F A3C Implementation
F.1. A3C Actions and State Representation
$\texttt{adopt}_{i}$: The action $\texttt{adopt}_{i}$, $i\in\{0,1,\ldots,k_{1}\}$ represents that the adversary abandons its adversarial fork and starts mining a new fork on top the longest public chain whose latest $i$ blocks are pruned. In the fixed block reward model, considering only the action $\texttt{adopt}_{0}$—namely adopting the longest chain—is enough as $\texttt{adopt}_{i}$ for $i>0$ cannot lead to a more profitable strategy. However, in the volatile block reward model, a deep block in the honest fork may have an extraordinary amount of transaction fees, incentivizing the adversary to orphan the block and steal its transaction fees. A block with an extraordinary fee can result from including a whale transaction as considered in (Bar-Zur et al., 2023) or from having a long generation time. Action $\texttt{adopt}_{i}$ is feasible if $i\leq l_{\mathcal{H}}$.
F.2. Objective Function
F.2.1. Before the difficulty adjustment
F.2.2. After the difficulty adjustment
Since the objective function defined in equation 22 is non-linear, we cannot naively assign the time step reward as is done in Section F.2.1. To address this challenge, we adopt the approach introduced in (Sapirshtein et al., 2017), which converts the non-linear optimization function into a linear one. We set the time step reward as follows:
where $\rho$ represents the time-averaged profit the adversary earns by following the optimal strategy, i.e., $\rho=O^{\mathcal{A}}(\pi^{*})$, where $\pi^{*}$ denotes the optimal strategy. At the start of training, the actual value of $\rho$ is unknown. Therefore, the training agents initialize with $\rho=O^{\mathcal{A}}(\pi^{\mathcal{H}})$, where $\pi^{\mathcal{H}}$ denotes the honest strategy. Certain test agents are tasked with measuring the time-averaged profit under the policy trained by the training agents at specific time steps during the training process. If this profit exceeds the highest time-averaged profit achieved so far, they prompt the training agents to update $\rho$ to the new highest time-averaged profit. For the detailed analysis of the non-linear objective function conversion readers are referred to (Sapirshtein et al., 2017). We briefly provide the intuition behind setting the time step reward as defined in equation 23 in the following. The value function of the initializing state $s_{0}$ under the time step reward introduced in equation 23 can be obtained as follows:
Another critical point is that the average time step duration after difficulty adjustment needs recalibration in the environment. In our implementation, the duration of each time step follows an exponential distribution with rate $\frac{1}{t^{B}}$. The duration of each time step is of great importance, as it affects the amount by which transaction weight increases in the mempool. Before the difficulty adjustment, $t^{B}$ is set to 10 minutes. However, after the difficulty adjustment, the value of block generation time $t^{B}$ needs to be adjusted according to the policy followed by the adversary. To calculate the block generation time after difficulty adjustment under policy $\pi$, we need to keep track of the ratio of the number of canonical blocks to the total number of blocks mined under policy $\pi$. To this end, the neural network outputs another value that aims to estimate the following sum:
F.3. Undercutting as a Continuous Action
In an A3C implementation, the neural network outputs a single scalar value representing the value of the current state, along with a probability distribution over the possible actions. For discrete actions, this probability distribution is a vector where each element represents the probability of selecting a specific action given the current state. For continuous actions, the network outputs the parameters of a probability distribution over the continuous action (Jia and Zhou, 2022). In our implementation, the neural network outputs a vector representing the probability distribution of the actions introduced in Section F.1. Additionally, it outputs two scalars representing the mean and variance of the duration of the undercutting action, which is modeled as a continuous action.
F.4. Implementation of the Undercutting Attack
F.5. Implementation Details
Appendix G Weight-Time Equation
Appendix H Selfish Mining After Difficulty Adjustment
As shown in Figure 15, under the mempool pattern observed on December 18, 2023, difficulty-adjusted selfish mining is more profitable in the fixed reward model compared to the volatile reward model. This suggests that, under selfish mining, the rate of increase in collected transaction fees is lower than the rate of increase in the adversarial block ratio. One of the reasons contributing to the higher profitability of selfish mining in the fixed reward model is the concavity of the time-fee function during this period. Figure 16 illustrates the time-fee function obtained under the mempool pattern from December 18, 2023, between 23:00 and 24:00 CET. To generate Figure 16, we simulated 100,000 blocks under this mempool pattern, collecting their generation times and fees. As depicted in Figure 16, the corresponding time-fee function for this period is concave.
Figure 16. The concave time-fee function.
To better understand the effect of the concavity of the time-fee function on selfish mining profitability, we review the following theorem.
Theorem H.1.